OpenStudy (lemanual):

Solve for x: -3|x + 7| = -12 x = 5 over 3 , x = − 19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (anonymous):

b

OpenStudy (texaschic101):

-3 | x + 7 | = - 12 --- divide both sides by -3 | x + 7 | = 4 x + 7 = 4 -(x + 7) = 4 x = 4 - 7 -x - 7 = 4 x = -3 -x = 4 + 7 -x = 11 x = -11 its B

OpenStudy (anonymous):

beaceuse -3|-3+7|=-12 -3|4| =-12

OpenStudy (lemanual):

Thanks guys

OpenStudy (texaschic101):

:)